OpenStudy (anonymous):

Factor the following expression completely: 45x^3y-80xy

OpenStudy (anonymous):

Hey Jim, any help?

OpenStudy (anonymous):

hey jim look, this is what I got: 45x^3y - 80xy = 5xy(9x^2 - 16) =5xy(3x+4)(3x-4)

OpenStudy (anonymous):

anyone?:(

OpenStudy (anonymous):

or is it =5xy(3x-4)(3x+4)?

OpenStudy (anonymous):

yes I did but I dont know if my answer is correct

OpenStudy (anonymous):

yea that looks good

OpenStudy (anonymous):

so is it 5xy(3x+4)(3x-4) or the other way around 5xy(3x-4)(3x+4)

OpenStudy (anonymous):

either of them are good it doesn't matter of the order your binomials are in

OpenStudy (anonymous):

oh okay thank you.

OpenStudy (anonymous):

No problem good luck

jimthompson5910 (jim_thompson5910):

sry was afk for a long while

OpenStudy (anonymous):

thanks, hey I have a word problem I need help with:( The area of a rectangular athletic field is represented by the expression 9y^2-25z^2 square meters. Write an algebraic expression to represent one possible set of dimensions (in the sense “length times width”) of the athletic field. Include correct units with your solution. Can anyone help me please? I need it explained to me so I'll know how to do it for finals!:(

OpenStudy (anonymous):

Anyone please

jimthompson5910 (jim_thompson5910):

using the difference of squares method, we can factor the given expression into (3y+5z)(3y-5z) So the dimensions are 3y+5z by 3y-5z Note: the difference of squares formula is a^2-b^2 = (a+b)(a-b) In this case, a = 3y and b = 5z